llp simplex method

Use theSimplex methodto solve the following problem:

Maximise  z= 4x1+ 6x2

Subject to:  

x1+x211

x1+x227

2x1+ 5x290

x1, x20.
Answer #1

While converting the inequality constraints to equality constraints and introducing slack variables:

-x1 + x2 +s1 = 11 (Since the inequality is ‘<=’, the slack variable needs to be added, to convert the inequality to equality)

x1 + x2 -s2 = 27 (Since the inequality is ‘>=’, the slack variable needs to be subtracted, to convert the inequality to equality)

The other constraint 2x1 + 5x2 = 90, does not require any slack variable since the constraint already has equality.

According to these constraints, the correct option is d. 2x1 + 5x2 = 90

The other constraints are incorrect since they do no match with the constraints above.


answered by: ANURANJAN SARSAM